40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 106 of 355  Not logged in ELOG logo
ID Date Author Type Category Subject
  12537   Fri Oct 7 10:29:57 2016 SteveUpdateVACRGA removed

The RGA is removed for repaire. It's volume at atmophere and sealed.. P4 reading of 38 Torr is not correct.

 

Attachment 1: RGAremoved.png
RGAremoved.png
  12536   Thu Oct 6 15:42:51 2016 LydiaUpdateSUSOutput matrix diagonalization

Summary: At the 40m meeting yesterday, Eric Q. gave the suggestion that we accept the input matrix weirdness and adjust the output matrix by driving each coil individually so that it refers to the same degrees of freedom. After testing this strategy, I don't think it will work. 

Yesterday evening I tested this idea by driving one ITMY coil at a time, and measuring the response of each of the free swing modes at the drive frequency. I followed more or less the same procedure as the standard diagonalization: responses to each of the possible stimuli are compared to build a matrix, which is inverted to describe the responses given the stimuli. For the input matrix, the sensor readings are the responses and the free swing peaks are the stimuli. For the output matrix, the sensors transformed by the diagonalized input matrix as the responses of the dofs which are compared, and the drive frequency peak associated with a coil output is the stimulus. However, the normalization still happens to each dof independently, not to each coil independently. 

The output matrix I got had good agreement with the ITMY input matrix in the previous elog: for each dof/osem the elements had the same sign in both input and output matrices, so there are no positive feedback loops. The relative magnitude of the elements also corresponded well within rows of the input matrix. So the input and output matrices, while radically different from the ideal, were consistent with each other and referred to the same dof basis. So, I applied these new matrices (both input and output) to the damping loops to test whether this approach would work. 

drive-generated output matrix: 

      UL      UR      LR       LL      SD
pit    1.701  -0.188  -2.000  -0.111   0.452  
yaw    0.219  -1.424   0.356   2.000   0.370  
pos    1.260   1.097   0.740   0.903  -0.763  
sid    0.348   0.511   0.416   0.252   1.000  
but    0.988  -1.052   0.978  -0.981   0.060

However, when Gautam attempted to lock the Y arm, we noticed that this change significantly impacted alignment. The alignment biases were adjusted accordingly and the arm was locking. But when the dither was run, the lock was consistently destroyed. This indicates that the dither alignment signals pass through the SUS screen output matrix. If the output matrix pitch and yaw columns refer instead to the free swing eigenmodes, anything that uses the output matrix and attempts to align pitch and yaw will fail. So, the ITMY matrices were restored to their previous values: a close to ideal input matrix and naive output matrix. We could try to change everything that is affected by the output matrices to be independent of a transformation to the free swing dof basis, and then implement this strategy. But to me, that seems like an unneccessary amount of changes with unpredictable consequences in order to fix something that isn't really broken. The damping works fine, maybe even better, when the input matrix is set by the output matrix: we define pitch, for example, to be "The mode of motion produced by a signal to the coils proportional to the pitch row of the naieve output matrix," and the same for the other dofs. Then you can drive one of these "idealized" dofs at a time and measure the sensor responses to find the input matrix. (That is how the input matrix currently in use for ITMY was found, and it seems to work well.) 

 

  12535   Thu Oct 6 03:56:43 2016 ericqUpdateLSCRevival Attempt

[ericq, Gautam, Lydia]

We spent some time tonight trying to revive the PRFPMI. (Why PR instead of DR? Not having to deal with SRM alignment and potentially get a better idea of our best-case PRG). After the usual set up and warm up, we found ourselves unable to hold on to the PRMI while the arms flash. In the past, this was generally solved through clever trigger matrix manipulations, but this didn't really work tonight. We will meditate on the solution.

  12534   Wed Oct 5 19:43:13 2016 gautamSummaryGeneralVent review

This elog is meant to review some of the important changes made during the vent this summer - please add to this if I've forgotten something important. I will be adding this to the wiki page for a more permanent record shortly.


Vent objectives:

  1. Clean ITMX, ITMY, ETMX, ETMY
  2. Replace ETMX suspension cage, replace Al wire standoffs with Ruby (sapphire?) standoffs.
  3. Shorten Y arm length by 20mm
  4. Replace 40mm aperture baffles in ETM chambers with 50mm black glass baffles

Optics, OSEM and suspension status:

ITMX & ITMY

  • ITMX and ITMY did not have any magnets broken off during the vent - all five OSEM coils for both were removed and the optic EQ stopped for F.C. cleaning.
  • Both HR and AR faces were F.Ced, ~20mm dia area cleaned.
  • The coils were re-inserted in an orientation as close to the original (as judged from photos), and the shadow sensor outputs were made as close to half their open values as possible, although in the process of aligning the arms, this may have changed
  • OSEM filter existense was checked (to be updated)
  • Shadow sensor open values were recorded (to be updated)
  • Checked that tables were level before closing up
  • The UL OSEM on ITMY was swapped for a short OSEM while investigating glitchy shadow sensor outputs. This made no difference. However, the original OSEM wasn't replaced. Short OSEM was used as we only had spare short OSEMs. Serial number (S/N 228) and open voltage value have been recorded, wiki page will be updated. Does this have something to do with the input matrix diagonalization weirdness we have been seeing recently?
  • ITMX seems to be prone to getting stuck recently, reason unknown although I did notice the LL OSEM was kind of close to the magnet while inserting (but this magnet is not the one getting stuck, as we can see this clearly on the camera - the prime suspect is UL I believe)
  • OL beam centering on in vacuum steering optics checked before closing up

ETMY

  • UL, UR and LR magents broke off at various points, and so have been reglued
  • No standoff replacement was done
  • Re-suspension was done using newly arrived SOS wire
  • Original OSEMs were inserted, orientations have changed somewhat from their previous configuration as we did considerable experimentation with the B-R peak minimization for this optic
  • OSEM filter status, shadow sensor open voltage values to be updated.
  • New wire suspension clamp made at machine shop is used, 5 in lb of torque used to tighten the clamp
  • HR face cleaned with F.C.
  • Optic + suspension towers air baked (separately) at 34C for curing of EP30
  • Checked that tables were level before closing up
  • 40mm O.D. black glass baffle replaced with 50mm O.D. baffle.
  • Suspension cage was moved towards ITMY by 19mm (measured using a metal spacer) by sliding along stop marking the position of the tower.

ETMX

  • Al wire standoffs <--> Ruby wire standoffs (this has changed the pitch frequency)
  • All magnets were knocked off at some point, but were successfully reglued
  • New SOS tower, new SOS wire, new wire clamp used
  • OSEM filter status, shadow sensor open voltage values to be updated.
  • OSEM orientation is close to horizontal for all 5 OSEMs
  • Table leveling was checked before closing up.
  • 40mm O.D. black glass baffle replaced with 50mm O.D. baffle.\

PRM

  • Some issues with the OSEMs were noticed, and were traced down to the Al foil caps covering the back of the (short) OSEMs, which are there to minimize the scattererd 1064nm light interfering with the shadow sensor, shorting one of the OSEMs
  • To mitigate this, all Al foil caps now have a thin piece of Kapton between foil and electrical contacts on rear of OSEM
  • No OSEMs were removed from the suspension cage during this process, we tried to be as gentle as possible and don't believe the shadow sensor values changed during this work, suggesting we didn't disturb the coils (PRM wasn't EQ stopped either)

SRM

  • The optic itself wasn't directly touched during the vent - but was EQ stopped as work was being done on ITMY
  • It initially was NOT EQ stopped, and the shift in table level caused by moving ITMY cage to the edge of the table for F.C. cleaning caused the optic to naturally drift onto the EQ stops, leading to some confusion as to what happened to the shadow sensor outputs
  • The problem was diagnosed and restoring ITMY to its original position made the OSEM signals come back to normal.

SR3

  • Was cleaned by drag wiping both front and back faces

SR2/PR2/PR3/BS/OMs

  • These optics were NOT intentionally touched during this vent
  • The alignment on the OMs was not checked before close-up
 

Other checks/changes

  • OL beams were checked on in-vacuum input and output steering mirrors to make sure none were close to clipping
  • Insides of viewport windows were checked for general cleanliness, given that we have found the outside of some of these to be rather dirty. Insides of viewports checked were deemed clean enough.
  • Steve has installed a new vacuum guage to provide a more realiable pressure readout. 
  • We forgot to investigate the weird behaviour of the AS beam that Yutaro and Koji identified in November. In any case, looks like the clipping of the AS beam is worse now. We will have to try and fix this using the PZT mounted OMs, and if not, we may have to consider venting again

Summary of characterization tasks to be done:

  1. Mode matching into the Y arm cavity given the arm length change
  2. HOM content in transmitted IR light from Y arm given the arm length change (Finesse models suggest that the 2f second order HOM resonance may have moved closer to the 00 resonance)
  3. Arm loss measurement
  4. Suspension diagonalization
  5. Check the Qs of the optics eigenmodes - should indicate if any of our magnets, reglued or otherwise, are a little loose
  12533   Wed Oct 5 19:10:04 2016 gautamUpdateGeneralArm loss measurement review

[ericq,gautam]

There are multiple methods by which the arm loss can be measured, including, but not limited to:

  1. Cavity ringdown measurement
  2. Monitoring IR arm transmission using ALS to scan the arm through multiple FSRs
  3. Monitoring the reflected light from the ITM with and without a cavity (Johannes has posted the algebra here)

We found that the second method is extremely sensitive to errors in the ITM transmissivity. The first method was not an option for a while because the AOM (which serves as a fast shutter to cut the light to the cavity and thereby allow measurement of the cavity ringdown) was not installed. Johannes and Shubham have re-installed this so we may want to consider this method.

Most of the recent efforts have relied on the 3rd method, which itself is susceptible to many problems. As Yutaro found, there is something weird going on with ASDC which makes it perhaps not so reliable a sensor for this measurement (unfortunately, no one remembered to follow up on this during the vent, something we may come to regret...sad). He performed some checks and found that for the Y arm, POY is a suitable alternative sensor. However, the whitening gain was at 0dB for the measurements that Johannes recently performed (Yutaro does not mention what whitening gain he used, but presumably it was not 0). As a result, the standard deviation during the 10s averaging was such that the locked and misaligned readings had their 'fuzz' overlapping significantly. The situation is worse for POX DC - today, Eric checked that the POX DC and POY DC channels are indeed reporting what they claim, but we found little to no change in the POX DC level while misaligning the ITM - even after cranking the whitening gain up to 40!

Eric then suggested deriving ASDC from the AS110 photodiode, where there is more light. This increased the SNR significantly - in a 10s averaging window, the fuzz is now about 10 ADC counts out of ~1500 (~<1%) as opposed to ~2counts out of 30 previously. We also set the gains of POX DC, POY DC and ASDC to 1 (they were 0.001,0.001 and 0.5 respectively, for reasons unknown).

I ran a quick measurement of the X arm loss with the new ASDC configuration, and got a number of 80 +/- 10 ppm (7 datapoints), which is wildly different from the ~250ppm number I got from last night's measurement with 70 datapoints. I was simultaneously recording the POX DC value, which yielded 40 +/- 10 ppm.

We also discovered another possible problem today - the spot on the AS camera has been looking rather square (clearly not round) since, I presume, closing up and realigning everything. By looking at the beam near the viewport on the AS table for various configurations of the ITM, we were able to confirm that whatever is causing this distortion is in the vacuum. By misaligning the ITM, we are able to recover a nice round spot on the AS camera. But after running the dither align script, we revert to this weirdly distorted state. While closing up, no checks were done to see how well centered we are on the OMs, and moreover, the DRMI has been locked since the vent I believe. It is not clear how much of an impact this will have on locking the IFO (we will know more after tonight). There is also the possibility of using the PZT mounted OMs to mitigate this problem, which would be ideal.


Long story short -

  1. Some more thought needs to be put into the arm loss measurement. If we are successful in locking the IFO, the PRG would be a good indicator of the average arm loss.
  2. There is some clipping, in vacuum, of the AS beam. It may be that we can fix this without venting, to be investigated.
 

GV Edit 8 Oct 2016: Going through some old elogs, I came across this useful reference for loss measurement. It doesn't talk about the reflection method (Method 3 in the list at the top of this elog), but suggests that cavity ringdown with the Trans PD yields the most precise numbers, and also allows for measuring TITM

  12532   Wed Oct 5 16:28:10 2016 gautamUpdateendtable upgradeEX laser power monitor PD installed

I installed a 10% BS to pick off some of the light going to the IR fiber, and have added a Thorlabs PDA55 PD to the EX table setup. The idea is to be able to monitor the power output of the EX NPRO over long time scales, and also to serve as an additional diagnostic tool for when ALS gets glitchy etc. There is about 0.4mW of IR power incident on the PD (as measured with the Ophir power meter), which translates to ~2500 ADC counts (~1.67V as measured with an Oscilloscope set to high impedance directly at the PD output). The output of the PD is presently going to Ch5 of the same board that receives the OL QPD voltages (which corresponds to ADC channel 28). Previously, I had borrowed the power and signal cables from the High-Gain Transmon PD to monitor this channel, but today I have laid out independent cabling and also restored the Transmon PD to its nominal state.

On the CDS side of things, I edited C1SCX to route the signal from ADC Ch28 to the ALS block. I also edited the ALS_END library part to have an additional input for the power monitor, to keep the naming conventions consistent. I have added a gain in the filter module to calibrate the readout into mW using these numbers. The channel is called C1:ALS-X_POWER_OUT, and is DQed for long-term trending purposes.

The main ALS screen is a bit cluttered so I have added this channel to the ALS overview MEDM screen for now..

  12531   Tue Oct 4 22:18:24 2016 JohannesUpdateGeneralX/YARM loss measurement

[gautam, johannes]

We let the PSL shutter closed overnight and observed the POXDC, POYDC and ASDC offsets. While POY has small fluctuations compared to the signal level, POX is worse off, and the drifts we observed live in the DC reading are in the same ballpark as the offset fluctuations. The POXDC level also unexpectedly increased suddenly without the PSL shutter being opened, which we can't explain. The data we took using POXDC cannot be trusted.

Even the ASDC occasionally shows some fluctuations, which is concerning because the change in value rivals the difference between locked and misaligned state. It turns out that the green shutters were left open, but that should not really affect the detectors in question.

We obtained loss numbers by measuring the arm reflections on the ASDC port instead. LSCoffsets was run before the data-taking run. For each arm we misaligned the respective other ITM to the point that moving it no longer had an impact on the ASDC reading. By taking a few quick data points we conclude the following numbers:

XARM: 247 ppm +/- 12 ppm
YARM: 285 ppm +/- 13 ppm

This is not in good agreement with the POYDC value. The script is currently running for the YARM for better statistics, which will take a couple hours.

ITMX is misaligned for the purpose of this measurement, with the original values saved.

GV edit 5Oct2016: Forgot to mention here that Johannes marked the spot positions on the ITMs and ETMs (as viewed on the QUAD in the control room) with a sharpie to reflect the current "well aligned" state.

Attachment 1: anomalousData.png
anomalousData.png
  12530   Tue Oct 4 11:02:31 2016 SteveUpdateVACRGA is out of order

The last good rga scan at vent 78  day 38

Quote:

 

Quote:

RGA background scan

Quote:

Vacuum Status: Chamber Open

All chamber annuloses are vented.  Vac Monitor screen is not communicating with gauges. The valve position indicator are working.

RGA is pumped by Maglev through VM2

 

 

 

 

Attachment 1: lastBGscan.png
lastBGscan.png
  12529   Tue Oct 4 02:59:48 2016 ericqUpdateGeneralmucking about

[ericq, gautam]

We poked around trying to figure out the X PDH situation. In brief, the glitchiness comes and goes, not sure what causes it. Tried temp servo on/off and flow bench fan on/off. Gautam placed a PD to pick off the pre-doubler AUX X IR light to see if there is some intermittant intensity fluctuation overnight. During non-glitchy times, ALSX noise profile doesn't look too crazy, but some new peak around 80Hz and somewhat elevated noise compared to historical levels above 100Hz. It's all coherent with the PDH control up there though, and still looks like smooth frequency noise...

NB: The IR intensity monitoring PD is temporarily using the high gain Transmon PD ADC channel, and is thus the source of the signal at C1:LSC-TRY_OUT_DQ. If you want to IR lock the X arm, you must change the transmon PD triggering to use the QPD.

Attachment 1: 2016-10-04_ALSXspectra.pdf
2016-10-04_ALSXspectra.pdf
  12528   Mon Oct 3 21:24:02 2016 JohannesUpdateGeneralXARM loss measurement

[gautam, johannes]

I started a script on Friday night to collect some data for a reflection armloss measurement of the XARM. Unfortunately there seemed to have been a hickup in some data transfer and some errors were produced, so we couldn't really trust the numbers.

Instead, we took a series of manual measurements today and made sure the interferometer is well behaved during the averaging process. I wrote up the math behind the measurement in the attached pdf.

The numbers we used for the calculations are the following:

While we average about 50 ppm +/-15 ppm for the XARM loss with a handful of samples, in a few instances the calculations actually yielded negative numbers, so there's a flaw in the way I'm collecting the data. There seems to be a ~3% drift in the signal level on the PO port on the order of minutes that does not show in the modecleaner transmission. The signals are somewhat small so we're closing the shutter over night to see if it could be an offset and will investigate further tomorrow. I went back and checked my data for the YARM, but that doesn't seem to be affected by it.

Attachment 1: ReflectionLoss.pdf
ReflectionLoss.pdf ReflectionLoss.pdf
  12527   Sat Oct 1 10:03:28 2016 ericqUpdateGeneralmucking about

Some things I did last night:

I measured the X PDH OLG, and turned the gain down by ~6dB to bring the UGF back to 10kHz, ~50deg phase margin, 10dB gain margin. However, the error signal on the oscilloscope remained pretty ratty. Zooming in, it was dominated by glitches occuring at 120Hz. I went to hook up the SR785 to the control signal monitor to see what the spectrum of these glitches looked like, but weirdly enough connecting the SR785's input made the glitches go away. In fact, with one end of a BNC connector plugged into a floating SR785 input, touching the other end's shield to any of the BNC shields on the uPDH chassis made the glitches go away.

This suggested some ground loop shenanigans to me; everything in the little green PDH shelves is plugged into a power strip which is itself plugged into a power strip at the X end electronics rack, behind all of the sorensens. I tried plugging the power strip into some different places (including over by the chamber where the laser and green refl PD are powered), but nothing made the glitches go away. In fact, it often resulted in being unable to lock the PDH loop for unknown reasons. This remains unsolved.

As Gautam and Johannes observed, the X green beat was puny. By hooking up a fast scope directly to the beat PD output, I was able to fine tune the alignment to get a 80mVpp beat, which I think is substaintially bigger than what we used to have. (Is this plus the PDH gain changed really attributable to arm loss reduction? Hm)

However, the DFD I and Q outputs have intermittent glitches that are big enough to saturate the ADC when the whitening filters are on, even with 0dB whitening gain, which makes it hard to see any real ALS noise above a few tens of Hz or so. Turning off the whitening and cranking up the whitening gain still shows a reasonably elevated spectrum from the glitches. (I left a DTT instance with a spectrum on in on the desktop, but forgot to export...) The glitches are not uniformly spaced at 120Hz as in the PDH error signal. However, the transmitted green power also showed intermittant quick drops. This also remains unsolved for the time being. 

  12526   Fri Sep 30 19:53:07 2016 gautamUpdateendtable upgradeX end IR pickoff fiber coupled

[johannes, gautam]

Today we re-installed the fiber coupler on the X-endtable to couple some of the PSL light into a fiber that runs to the PSL table, where it is combined with a similar PSL pickoff to make an IR beat between the EX AUX laser and the PSL. The main motivation behind this was to make the process of finding the green beatnote easier. We used JAMMT (just another mode matching tool) to calculate a two lens solution to couple the light into the collimator - we use a +200mm and -200mm lens, I will upload a more detailed mode matching calculation + plot + picture soon. We wanted to have a beam waist of 350um at the collimator, a number calculated using the following formula from the Thorlabs website:

d =4\lambda (\frac{f}{\pi*MFD})

where d is the diameter of the output beam from the collimator, f is the collimating lens focal length and MFD is 6.6um for the fiber we use.

There is ~26mW of IR light coming through the BS after the EX AUX - after playing around with the 6 axis stage that the coupler is mounted on, Johannes got the IR transmission to the PSL table up to ~11.7mW. The mode matching efficiency of 45% is certainly not stellar, but we were more curious to find a beat and possibly measure the X arm loss so we decided to accept this for now - we could probably improve this by moving the lenses around. We then attenuated the input beam to the fiber by means of an ND filter such that the light incident on the coupler is now ~1.3mW, and the light arriving at the PSL table from the EX laser is ~550uW. Along with the PSL light, after the various couplers, we have ~500uW of light going to the IR beat PD - well below its 2mW threshold.

The IR beat was easily found with the frequency counter setup. However, there was no evidence of a green beat. So we went to the PSL table and did the near-field-far-field alignment onto the beat PD. After doing this, we were able to see a beat - but the amplitude was puny (~-60dBm, we are more used to seeing ~-20dBm on the network analyzer in the control room). Perhaps this can be improved by tweaking the alignment onto the PD while monitoring the RF output with an oscilloscope.

Moreover, the green PDH problems with the X end persist - even though the arm readily locks to a TEM00 mode, it frequently spontaneously drops lock. I twiddled around with the gain on the uPDH box while looking at the error signal while locked on a oscilloscope, but was unable to mitigate the situation. Perhaps the loop shape needs to be measured and that should tell us if the gain is too low or high. But ALS is getting closer to the nominal state...

Johannes is running his loss measurement script on the X arm - but this should be done by ~10pm tonight.

 

  12525   Fri Sep 30 10:37:57 2016 SteveUpdateVACvacuum VME machines rebooted

[ Gautam and Steve ]

c1vac1 and c1vac2 were rebooted and the gauges are communicating now. V1, VA6, V5 and V4 were closed and disconnected to avoid unexpected valve switching. All went smoothly.

The new ITcc gauge is at 1e-5 Torr as CC1   This is the gauge that should be logged in slow channel.

TP2 fore line dry pump was replaced this morning after 382 day of operation.

TP3 dry pump is very noisy, but it's pressure still 47 mTorr

Quote:

[ericq, Gautam, Steve]

Following roughly the same procedure as ELOG 11354, c1vac1 and c1vac2 were rebooted. The symptoms were identical to the situation in that ELOG; c1vac1 could be pinged and telneted to, but c1vac2 was totally unresponsive. 

The only change in the linked procedure was that we did not shut down the maglev. Since I unwittingly had it running for days without V4 open while Steve was away, we now know that it can handle shorter periods of time than that...

Upon reboot, many channels were readable again, unfortunately the channels for TP2 and TP3 are still blank. We were able to return to "Vacuum normal state," but because of unknowned communication problems with VM1's interlock, we can't open VM1 for the RGA. Instead we opened VM2 to expose the RGA to the main IFO volumn, but this isn't part of the "Normal" state definite, so things currently read "Undefined state".

 

Attachment 1: rebooted.png
rebooted.png
  12524   Thu Sep 29 20:21:29 2016 JohannesUpdateGeneralYARM loss measurement

[Gautam, Johannes]

I scripted a series of YARM DC reflectivity measurements last night alternating between locked state and unlocked state (with ETMY misaligned) for measuring the after-vent armloss. The general procedure is based on elog 11810, but I'll also give a brief summary here.

  • To measure the locked reflectivity the dither script is executed with a stop condition that depends on the rms values of its error signals.
  • The dithering is stopped, and while the arm is locked the reflected power is recorded from both POX/POY DC and ASDC, as well as the mode cleaner transmission for normalization.
  • The arm locking is switched off, and ETMY moved to is 'misaligned' position. This gets rid off unwanted mode flashes.
  • In the unlocked state the same quantities are recorded.
  • Rinse and repeat for a set number of times (for this run I set it to 100 and left the interferometer alone).

I did this back in June (but strangely never posted what I found, shame on me). What I found back then was a YARM loss of 237 ppm +/- 41 ppm and an XARM loss of 501 ppm +/- 105 ppm

Last night's data indicates a YARM loss of 143 ppm +/- 24 ppm after cleaning with first contact. yeslaugh

THIS IS STILL ASSUMING THAT THE MODE-MATCHING HASN'T CHANGED. We had however moved ETMY closer to ITMY during the vent by 19mm. Gautam and I had some trouble setting up the ALS to confirm the mode-matching, but we're in the process of recovering the XARM IR beat.

  12523   Thu Sep 29 16:19:29 2016 LydiaUpdateSUSFree swing eigenmodes

[Lydia, Teng]

Motivated by the strange pitch/yaw coupling behavior we ran into while doing diagonalization, we looked at the oplev pitch and yaw free swing spectra for all 4 test masses (see attachment 1). We saw the same behavior there: At the peak frequencies for the angular degress of freedom, the oplevs saw significant contributions from both pitch and yaw. We also examined the phase between pitch and yaw at these peaks and found that consistently, pitch and yaw were in phase at one of the resonance frequencies and out of phase at the other (ignoring the pos and side peaks). 

This corresponds physically to angular motion about some axis that is diagonal, ie not perfectly vertical or horizontal. If we trust the oplev calibration, and Eric says that we do, then the angle of this axis of rotation with the horizontal (pitch axis) is

 \theta = \arctan \frac{Y\left ( f_{peak} \right )}{P\left ( f_{peak} \right )}  

Where Y and P are yaw and pitch ASD values. This will always give an angle between 0 and 90 degrees; which quadrant the axis of rotation occupies can be dermined by looking at the phase between pitch and yaw at the same frequencies. 0 phase means that the axis of rotation lies somewhere less than 90 degrees counterclockwise from the horizontal as viewed from the AR face of the optic, and a phase of 180 degrees means the axis is clockwise from horizontal (see attachment 2). Qualitatively, these features show up the same way for segments of data taken at different times. In order to get some quantitative sense of the error in these angles, we found them using spectrogram values with a bandwidth of 0.02 Hz averaged over 4000 seconds. 

Results (all numbers in degrees unless otherwise specified):

ITMY
peak 1 ( 0.692  Hz):
mean: 24.991
std: 1.23576
ptich/yaw phase: -179.181
peak 2 ( 0.736  Hz):
mean: 21.7593
std: 0.575193
pitch/yaw phase: 0.0123677

 

ITMX
peak 1 ( 0.502  Hz):
mean: 17.4542
std: 0.745867
ptich/yaw phase: -179.471
peak 2 ( 0.688  Hz):
mean: 74.822
std: 0.455678
pitch/yaw phase: -0.43991

 

ETMX
peak 1 ( 0.73  Hz):
mean: 43.1952
std: 1.54336
ptich/yaw phase: -0.227034
peak 2 ( 0.85  Hz):
mean: 60.7117
std: 0.29474
pitch/yaw phase: -179.856

ETMY
peak 1 ( 0.724  Hz):
mean: 78.2868
std: 2.18966
ptich/yaw phase: 6.03312
peak 2 ( 0.844  Hz):
mean: 26.0415
std: 2.10249
pitch/yaw phase: -176.838

ETMY and ITMX both show a more significant (~4x) contribution from pitch on one peak, and from yaw on the other. This is reflected in the fact that they each have one angle somewhat close to 0 (below 30 degrees) and one close to 90 (above 60 degrees). The other two test masses don't follow this rule, meaning that the 2 angular frequency peaks do not correspond to pitch and yaw straightforwardly. 

Also, besides ITMX, the axes of rotation are at least several degrees away from being perpendicular to each other. 

 

Attachment 1: 05.png
05.png
Attachment 2: SUS_eigenmodes.png
SUS_eigenmodes.png
  12522   Thu Sep 29 09:49:53 2016 ranaUpdateGeneralmucking about

With the WFS and OL, we never have figured out a good way to separate pit and yaw. Need to figure out a reference for up/down and then align everything to it: quad matrix + SUS output matrix

  12521   Wed Sep 28 04:27:33 2016 ericqUpdateGeneralmucking about

PMC was terribly misaligned. The PMCR camera seems to have drifted somewhat off target too, but I didn't touch it.

Realigned ITMX for the nth time today.

Finding ALSY beatnote was easy, ALSX eludes me. I did a rough one-point realignment on the X beat PD which is usually enough, but it's probably been long enough that near/far field alignmnet is neccesary. 

ALSY noise is mostly nominal, but there is a large 3Hz peak that is visible in the spot motion, and also modulates the beat amplitude by multiple dBs.

It looked to me that the ETMY oplev spot was moving too much, which led me to measure the oplev OLGs. There is some wierd inter-loop interference going on between OLPIT and OLYAW. With both on (whether OSEM damping is on or off, so input matrix shenanigans can't be to blame) there is a very shallow "notch" at around 4.5Hz, which leads to very little phase at 3Hz, and thus tons of control noise. Turning the OL loop not being measured off makes this dip go away, but the overall phase is still signfinicantly less than we should have. I'm not sure why. I'll just show the PIT plot, but things look pretty much the same for YAW. 


I did some more ETMX tests. Locked arm, raised the servo output limit to 15k, then increased the gain to make the loop unstable. I saw the SUS LSC signals go up to tens of thousands of counts when the unlock happened. I did this a dozen times or so, and every time the ETM settled in the same angular position according to the oplev.

Right now, another hysteresis script is running, misaliging in pitch and yaw. Amplitude 1V in each direction. So far, everything is stable after three on/off cycles.

Attachment 1: alscheck.pdf
alscheck.pdf
Attachment 2: weird_olpit.pdf
weird_olpit.pdf
  12520   Tue Sep 27 18:04:50 2016 LydiaUpdateSUSITMX slow channels down, ITMY diagonalization update

[Teng, Lydia]

When we plugged in the back cables yesterday on the whitening boxes after switching them, two of the ITMX PDMon channels (UR and LR) got stuck at 0. This caused me to believe ITMX was still stuck even when it was freed. However, it was left in a stuck state overnight and freed again today after this issue was discovered. The alignment sliders have been set to 0 as a safety net to keep ITMX from getting stuck again if c1susaux is restarted again. We switched the cables back and the problem was still there.

The ITMY UL whitening filter problem, which the cables were originally switched to diagnose, was also still there. Ericq suggested we turn off all the whitening filters in order to get diagonalization data that would not show a phase difference between coils. We ran the diagonalization again with all the dewhitening filters off and got much cleaner results, with no visible cross-coupling peaks remaining between the degrees of freedom (see attachemnt 1). We did not apply this matrix to the damping, however, because there are elements which have the wrong sign compared to the ideal matrix. Significant adjustments to the output matrix will probably need to be made if this result is to be used. We also verified that the phase problem had been solved in DTT, where we saw the same sign discrepancies as in the matrix below. 

Damping can be turned back on, using the old, non-diagonalized matrix currently in effect. There is enough free swing data to diagonalize ITMY now, so feel free to mess with it. 

Matrix (wrong signs red, suspiciously small elements orange):

           pit     yaw     pos         side    butt
UL    1.633   0.138   1.224   0.136   0.984  
UR   -0.202  -1.768   1.179   0.132  -1.028  
LR   -2.000   0.094   0.776   0.107   1.001  
LL   -0.165   2.000   0.821   0.111  -0.987  
SD    0.900   1.131  -1.708   1.000  -0.107  

 

Attachment 1: ITMY_diagsuccess.pdf
ITMY_diagsuccess.pdf
  12519   Tue Sep 27 08:49:47 2016 SteveUpdateSUSseismic activity is up

The earth quake shook ITMX free for a  short while.

 

Attachment 1: 4.3mSaltonSee.png
4.3mSaltonSee.png
Attachment 2: ITMXstuck.png
ITMXstuck.png
  12518   Mon Sep 26 19:48:09 2016 LydiaUpdateSUSITMX stuck again, ITMY whitening issue

This afternoon around 2:45, ITMX started ringing up at ~.9Hz for about a minute and then got stuck again. When I noticed this evening, I tried to free it with the alignment sliders but was unable to see any signal on UL or UR. It also looks like the damping for ITMY was turned off at the same time ITMX got stuck (not at the start of its ring up). SRM also has a spike in its motion at this time, and another one minute later that ended up with the LR OSEM at a much higher level, though the mirror does not appear to be stuck. We didn't see any strange behavior from any of the other optics.

Teng and I were working on diagnosing a problem with the ITMY UL whitening, but by the time we disconnected any applicable cables, the damping for ITMY was already off. Later we unplugged the ITMX PD whitening cables after verifying that the ITMX damping was also already off. This problem may have occured earlier, while Teng, Eric, and I were examining and pushing in the cables at 1X5 without unplugging anything.

We found that the reason for the bad phase on the ITMY free swing data is because the whitening filter for UL is not being properly turned on. We are in the process of investigating the source of this problem. Right now all the cables to the PD whitening boxes in 1X5 are switched between ITMY and ITMX.
 

Attachment 1: 44.png
44.png
Attachment 2: 26.png
26.png
  12517   Sat Sep 24 11:04:00 2016 ericqUpdateSUSETMX hysteresis test

Seems like the angular position was fairly stable, though there is some change in the ETMX pitch that could be hysterisis or normal drift. I didn't mention it explicity in the previous log, but the misalignment was purely in pitch. I'll give it another shot with a bigger misalginment, and maybe a mix of pitch and yaw.

  12516   Fri Sep 23 01:09:04 2016 ericqUpdateSUSETMX hysteresis test

I had hoped to do some ALS work, but I realized too late that we loaned our HP analyzer to Andrew. I decided instead to do some ETMX testing. 

I have a script running that'll misalign both ETMs and back by about 0.5mrad with half hour rests in between. It'll be done around 6AM.

  12515   Thu Sep 22 22:52:08 2016 ericqUpdateGeneralDamping found to be on

Just a heads up, it looks like the damping came on at around 8:30pm. Not sure why. 

  12514   Thu Sep 22 20:18:27 2016 LydiaUpdateGeneralAcromag Progress

We moved the Acromag and its power supply to the X end, where we connected it to the diagnostic output of the NPRO controller. We renamed the channels to be descriptive of the pin outputs as described in the laser manual. We were able to recover readouts similar to those we found with a multimeter. 

We should figure out how to set up the channels on the front end machines: right now they are accessed through a tmux session running on pianosa. Once we are confident in the operation, we will make a box to contain the Acromag and wire connections and move the setup to connect to the PSL controller. 

  12513   Thu Sep 22 20:01:47 2016 LydiaUpdateGeneralITMX freed, all optics kicked

Rana came by and freed ITMX again. I think it shouldn't be a problem for me to free it if it happens again. 

In hopes of getting better SNR on the free swing spectra, we kicked all optics at around 7pm. The damping should come back on a little after midnight. ITMX did not get stuck after this kick. 

  12512   Thu Sep 22 15:44:21 2016 SteveUpdateGeneralITMX magnets are stucked again

 

 

Attachment 1: StuckAgain.png
StuckAgain.png
  12511   Wed Sep 21 09:04:57 2016 SteveUpdateGeneral8 hours recovery progress

Good 8 hours

Quote:

The misalignment wasn't as bad as I had intially feared; the spot was indeed pretty high on ETMX at first. Both transmon QPDs did need a reasonable amount of steering to center once the dither had centered the beam spots on the optics.

Arms, PRMI and DRMI have all been locked and dither aligned. All oplevs and transmon QPDs have been centered. All AS and REFL photodiodes have been centered. 

Green TM00 modes are seen in each arm; I'll do ALS recovery tomorrow. 

 

Attachment 1: 8hrs.png
8hrs.png
  12510   Wed Sep 21 01:08:02 2016 ericqUpdateGeneralFurther recovery progress

The misalignment wasn't as bad as I had intially feared; the spot was indeed pretty high on ETMX at first. Both transmon QPDs did need a reasonable amount of steering to center once the dither had centered the beam spots on the optics.

Arms, PRMI and DRMI have all been locked and dither aligned. All oplevs and transmon QPDs have been centered. All AS and REFL photodiodes have been centered. 

Green TM00 modes are seen in each arm; I'll do ALS recovery tomorrow. 

  12509   Tue Sep 20 17:04:46 2016 SteveUpdateElectronicsREF33

REF33 was removed for taking picture of the bare C30362 InGaAs photodiode per Rana's request. All other rf photodiodes have their glass cover on.

Note: it is back to it's place but this pd will need alignment!

The small steering mirror was completly lose before it was removed.

Attachment 1: A005_-_20160920_135529_-_Shortcut.lnk.bmp
  12508   Tue Sep 20 10:45:06 2016 ranaUpdateGeneralFurther recovery progress

Rana suspicious. We had arms locked before pumpdown with beams on Transmon PDs. If they're off now, must be beams are far off on the mirrors. Try A2L to estimate spot positions before walkin the beams too far.

  12507   Mon Sep 19 22:03:10 2016 ericqUpdateGeneralFurther recovery progress

[ericq, Lydia, Teng]

Brief summary of this afternoon's activities:

  • PMC alignment adjusted (Transmission of 0.74)
  • IMC locked, hand aligned. Tranmission slightly over 15k. Measured spot positions to be all under 2mm.
  • Set DC offsets of MC2 Trans + WFS1 + WFS2 (WFS2 DC offsets had wandered so much that DC "centered" left some quadrants almost totally dark)
  • Set demod offsets of WFS1+WFS2
    • Note to self: WFS script area is a mess. I can never remember which scripts are the right ones to run. I should clean this up
  • WFS loops activated, tested. All clear.
  • Locked Yarm, dither aligned. Transmission 0.8
  • Moved BS to center ITMY reflection on AS camera
  • Misaligned ETMY, aligned PRM to make a flashing PRY AS beam. REFL camera spot confirmed to be on the screen, which is nice
  • Wandered ITMX around until its AS spot was found. ITMX OSEMs not too far from their half max. (todo: update with numbers)
  • Wandered SRM around until full DRMI flashes seen
  • Centered all vertex oplevs
  • Made a brief attempt at locking X arm, could only get some crazy high order mode to lock. BS and ITMX alignments have changed substantially from the in-air locks, so probably need to adjust ETMX much more.

Addendum: I had a suspicion that the alignment had moved so much, we were missing the TRX PDs. I misaligned the Y arm, and used AS110 as a proxy for X arm power, as we've done in the past for this kind of thing. Indeed, I could maximize the signal and lock a TM00 mode. Both the high gain PD and QPD in the TRX path are totally dark. This needs realignment on the end table.

  12506   Mon Sep 19 13:57:21 2016 ranaUpdateGeneralnever post EPS files in the ELOG. Ever.

http://tex.stackexchange.com/questions/2092/which-figure-type-to-use-pdf-or-eps

  12505   Mon Sep 19 13:25:03 2016 TengUpdateElectronicsSatellite Amplifier

 

In order to figure out the difference betweent simulated result and measurement, I tried to measuren the electronic noise by following ways as show in attachment 1

1.measure from the satellite box by SR785 at ETMY ,calibrate to counts by divide by 3267.8. while at that conditin, the set up is in suspension.

2. measure after ADC by diagnostics test tools, with set up on table in history and on uspension currently.

3. use the caculated butterfly channel.

the results are shown in attachmemt 2. The overall nosie level are still much higher than simulation.

 

 

Quote:

 

If we have some data with one of the optics clamped and the open light hitting the PD, or with the OSEMs removed and sitting on the table, that would be useful for evaluating the end-to-end noise of the OSEM circuit. It seems like we probably have that due to the vent work, so please post the times here if you have them.

The ETMX OSEMs have been attached to its Satellite box and plugged in for the last 10 days or so, with the PD exposed to the unobstructed LED. I pulled the spectrum of one of the sensors (mean detrended, I assume this takes care of removing the DC value?). The DQed channels claim to record um (the raw ADC counts are multiplied by a conversion factor of 0.36). For comparison, re-converted the y-axis for the measured curve to counts, and multiplied the total noise curve from the LISO simulation by a factor of 3267.8cts/V (2^16cts/20V) so the Y axis is noise in units of counts/rtHz. At 1Hz, there is more than an order of magnitude difference between the simulation and the measurement which makes me suspect my y-axis conversion, but I think I've done this correctly. Can such a large discrepancy be solely due to thick film resistors?

 

  12504   Mon Sep 19 11:11:43 2016 ericqUpdateSUSAlignment status

[ericq, Steve]

ITMX is free, OSEM signals all rougly centered. 


This was accomplished by rocking the static alignment (i.e. slow controls) pitch and yaw offsets until the optic broke free. This took a few volts back and forth. At this point, I tried to find a point where the optic seemed to freely swing, and hopefully have signals in all 5 OSEMS. It seemed to be free sometimes but mostly settling into two different stationary states. I realized that it was becoming torqued enough in pitch to be leaning on the top-front or top-back EQ stops. So, I slowly adjusted the pitch from one of these states until it seemed to be swinging a bit on the camera, and three OSEM signals were showing real motion. Then, I slowly adjusted the pitch and yaw alignments to get all OSEMS signals roughly centered at half of their max voltage.

  12503   Sun Sep 18 16:18:05 2016 ranaUpdateSUSAlignment status

susaux is responsible for turning on/off the inputs to the coil driver, but not the actual damping loops. So rebooting susaux only does the same as turning the watchdogs on/off so it shouldn't be a big issue.

Both before and after would be good. We want to see how much bias and how much voltage from the front ends were applied. l1susaux could have put in a huge bias, but NOT a huge force from the damping loops. But I've never seen it put in a huge bias and there's no way to prevent this anyway without disconnecting cables.

I think its much more likely that its a little stuck due to static charge on the rubber EQ stop tips and that we can shake it lose with the damping loops.

  12502   Sat Sep 17 16:51:01 2016 LydiaUpdateSUSAlignment status

Here's the timeseries plots. I've zoomed in to right after the problem- did you want before? We pretty much know what happened: c1susaux was restarted from the crate but the damping was on, so as soon as the machine came back online the damping loops sent a huge signal to the coils. (Also, it seems to be down again. Now we know what to do first before keying the crate.) It seems like both right side magnets are stuck, and this could probably be fixed by moving the yaw slider. Steve advised that we wait for an experienced hand to do so. 

Quote:

All is not lost. I've stuck and unstuck optics around a half dozen times. Can you please post the zoomed in time series (not trend) from around the time it got stuck? Sometimes the bias sliders have to be toggles to make the bias correct. From the OSEM trend it seems like it got a large Yaw bias. May also try to reseat the satellite box cables and the cable from the coil driver to the cable breakout board in the back of the rack.

 

Attachment 1: Screenshot_from_2016-09-17_16-45-00.png
Screenshot_from_2016-09-17_16-45-00.png
  12501   Sat Sep 17 02:00:23 2016 ranaUpdateSUSAlignment status

All is not lost. I've stuck and unstuck optics around a half dozen times. Can you please post the zoomed in time series (not trend) from around the time it got stuck? Sometimes the bias sliders have to be toggles to make the bias correct. From the OSEM trend it seems like it got a large Yaw bias. May also try to reseat the satellite box cables and the cable from the coil driver to the cable breakout board in the back of the rack.

  12500   Fri Sep 16 19:48:52 2016 LydiaUpdateGeneralAlignment status

Today the Y arm was locking fine. The alignment had drifted somewhat so I ran the dither and TRY returned to ~0.8. However, the mode cleaner has been somewhat unstable. It locked many times but usually for only a few minutes. Maybe the alignment or autolocker needs to be adjusted, but I didn't change anything other than playing with the gain sliders (which didn't seem to make it either better or worse).

ITMX is still stuck.

  12499   Fri Sep 16 19:14:27 2016 LydiaUpdateSUSDiagonalization

[Lydia, Teng]

We built matrices for ITMY and ETMY by driving one degree of freedom at a time with awggui, while the damping was on. These have been applied to the damping loops.

  • Each segment of data is 1000s long and each dof was driven at 0.25 Hz.
  • These matrices are much closer to the ideal matrix and have no wrong signs. We believe they represent the relative sensitivity of the OSEMs to the degrees of freedom much more accurately. This is because the free swinging modes are not actually pitch, yaw, etc, but some linear combination of these. However, the damping actuates on pitch, yaw, etc. So we should isolate the degrees of freedom by driving them one at a time instead of just looking at free swinging peaks.
    • Attachment 1: An example of the dof spectra, calculated using the default input matrix, when ETMY YAW was driven at 0.25 Hz.
    • Attachment 2: The same OSEM sensor data, with the dofs calculated using the matrix found from this data. There is still a significant peak in pitch, but the other dofs are significantly suppressed.
    • Attahcment 3: The same data again, but the dofs are measured with the input matrix calculated by the free swinging data. This achieves much less suppression than the new matrix. Obviously this is not exactly a fair comparison because the new matrix was generated with this data, but the method of measuring OSEM responses by driving peaks has a much close relationship between what it measured (the OSEM response), and how the matrix is used (by damping loops which drive the coils in much the same way as awggui).
  • The phase problems seem to be mostly solved. Both Y arm test masses have some phase warnings, but they mostly occur with side. This can happen because the ideal matrix elements are 0, so the real parts are small. If there is no strong coupling then there is no reason to expect the background spectrum to be in phase with the peak. Other phase differences are small; most less than 5 degrees, a couple between 5 and 10 degrees. This may still merit further investiagtion.
  • Comparing the damping results for ITMY with the old (based on free swinging data) and new (based on driven data), we see the 1Hz peak suppressed by ~35% and the noise above 1Hz generally suppressed by ~25-30% . There is, however, significantly more movement between 0.5 and 1 Hz, maybe because the fundamental physical modes are not being directly measured and suppressed. Overall this seems like an improvement.

GPS times:

ITMY

Pitch:1158085097 Yaw: 1158086537 Pos: 1158089237 Side: 1158087977

ETMY

Pitch: 1158095897 Yaw: 1158097577 Pos: 1158099377 Side: 1158100817

Attachment 1: ETMY_yawdrivedefault.png
ETMY_yawdrivedefault.png
Attachment 2: ETMY_yawdrivenew.png
ETMY_yawdrivenew.png
Attachment 3: ETMY_yawdriveold.png
ETMY_yawdriveold.png
Attachment 4: 57.png
57.png
  12498   Fri Sep 16 14:15:28 2016 steveUpdatePEMpartical counts

South end flow bench and both clean room assembly flow benches measured zero counts for 0.3 and 0.5 micron size particales.

The counting efficiency of 0.5 micron is 100%

 

  location

 

 

        0.3 micron particles / cf min

 

 

      0.5 micron particles / cf min

 

 

 counter

 MET

 ONE

 

       effect
 ITMY table  35,000  3,000   #3  
 ETMY table  35,000  3,000   #3  
 ITMX table  24,000  2,400   #3  PSL HEPA at 50V
 ITMX wall mounted  -  2,500   #1  
 ETMX work bench    5,000     600   #3   flow bench on  

 

The PSL HEPA performance was measured at the center of the table with MET ONE #3

 Voltage Variac particles /cf min   particles /cf min
 particle size  0.3 micron  0.5 micron
 60  0  0
 50  10  0
 40  1,400  110
 20  1,400  100
Attachment 1: 80days.png
80days.png
  12497   Thu Sep 15 18:37:20 2016 LydiaUpdateSUSDiagonalization

[Teng, Lydia]

  • We fixed the 60Hz filter on ITMY. This improved the phase problems somewhat but one coil (UL) is still about 12 degrees out of phase compared to the others for all the dofs. Is there some other place where a filter coule be applied to just one coil sensor? I pressed the "Load coefficients" button for UL, so maybe that will have helped.
  • We want to interpret the coil signals to have an accurate measurement of each dof. This means what the input matrix should describe is the dependence of each dof on the OSEM signals, which is found by inverting the matrix which describes the sensitivity of each OSEM to changes in that degree of freedom.
    • We looked at the spectra of the individual coils for ITMY and ETMY (See attachment 1 & 2). The coupling between some coils and applicable resonance peaks is very weak (~0.1 times the sensitivity of the other coils).
    • However, when a certain degree of freedom, e.g. pitch, is deliberately driven using awggui, the response of the ITMY coils is clear on the StripTool and is about the same magnitude for all of the face OSEMS. So, it seems like the diagonalization script does not always succeed at measuring the relative sensitivity of the OSEMs to the degrees of freedom.
    • This may be because the fundamental swing modes experienced by the free swinging pendulum are not the same as what we measure as pitch, yaw, etc. This could be possible if the wire tension is not the same on both sides. For ITMY, the spectra imply that the funamdental frequencies are actually at some linear combinations of pitch and yaw, swinging about a diagonal axis that results in a much weaker response for some of the OSEMS. Calling these peaks pitch and yaw may be inaccurate. Certainly they do not indicate the true relative sensitivity of the coils.
    • We propose an alternate approach to measuring this sensitivity: drive one dof at a time with awggui, take a spectrum (less resolution is ok because we already know the drive frequency), and measure the sensing matrix values for that dof the same way as before, but using a spectral peak that decribes motion that we know is purely pitch. Repeat this for all 4 dofs that we can actuate on, then compile these results into a sensing matrix and take the inverse.
Attachment 1: ETMY_osemspec.png
ETMY_osemspec.png
Attachment 2: ITMY_osemspec.png
ITMY_osemspec.png
  12496   Thu Sep 15 10:36:32 2016 SteveUpdateVACpumpdown 79 day 3

IFO pressure 3.7e-5 Torr at new cold cathode InstruTech - Hornet

New items in vacuum:

1, ETMX sus tower with new baked sus wire, EP30-2 epoxied magnets, same at different locations also........ ..........  and 2 ruby wirestandoffs.

2, First Contact cleaned arm test masses only. This technic was a 1st time use in our vacuum system.

3, 50 mm ID green glass baffles at the ends

4, witness mirrors at ETMX and ITMY (old oplev mirrors) We observed a very dusty system: sides of optics, towers and tables were wiped. Hepa tents used at Y arm and BS

5, new pirani, cc gauge and 1.5" right angle valve

 

Quote:

The pumpdown started at 9-12-2016

The IFO is at 5e-5 Torr vacuum normal after 73 day at atm.

ITMX needed to be freed and ETMY-UL is still misbehaving occasionally.

New pirani   and cold cathode gauges added at this vent. They were baked at 100 C for 6 hrs under vacuum.

Go to the Vac Rack to read IFO pressure from the gauge itself when Vac. Monitor is blank as it is now !

 

Quote:

P1 IFO pressure is 1 mTorr, valve configuration: vacuum normal, annulosses are pumped, RGA is not pumped.

THANKS to Chris !

The shutters can be opened with high power.

No communication error message still exist.

I will reboot c1vac1 and c1vac2 to get gauge communication with medm tomorrow.

 

 

 

 

Attachment 1: vent78for73d.png
vent78for73d.png
  12495   Wed Sep 14 20:27:03 2016 LydiaUpdateSUSDiagonalization

Today the main optics were free swinging for several hours, so I attempted diagonalization in vacuum.

  • ITMY still has bad phases. I looked at the spectra for this and other optics, and it looks like the other optics have the 60Hz line notched out for all coils while ITMY only has it notched on the side coil. (Using C1:SUS-ITMY_SENSOR channels). Where is this controlled from, and could it be the source of the issue? 
    • I tried using a different coil as the "standard," with the other coils compared against it in tfestimate. Default is UL, I tried UR and LL. The phase problems were still present for ITMY, but the script was still working fine for other optics.
    • The phase difference between coils is different for different start times.
    • A short segment of the time series for ITMY shows significantly more high frequency noise than for other optics at the same time.
  • The ETMY matrix for vacuum has the wrong sign for UL coupling to pitch! The diagonalization results look OK on the graph, but the butterfly mode still has small peaks (See attachment 1). When the individual coil spectra are plotted, the angular degrees of freedom show very weak coupling for UL to pitch, and LL to yaw. We initially replaced the matrix on the MEDM screen with the one generated by the script. After realizing this, the PIT row was changed to 1 1 -1 -1 0, but the effectiveness of the damping on the locked transmission fluctuations was about the same both ways.
Attachment 1: ETMY_diag.png
ETMY_diag.png
  12494   Wed Sep 14 20:05:32 2016 LydiaUpdateGeneralITMX magnets are stucked

When I restarted c1susaux yesterday, I didn't know that I needed to disable the coil outputs first. So when it came back online, it attempted to damp all the vertex area optics and ITMX got stuck frown

We should make a note in the Computer Restart Procedures wiki page indicating the importance of disabling the coils before rebooting c1susaux, c1auxex, and c1auxey. Today c1auxey was rebooted properly without incident. If the slider values etc go back to their previous values on their own, is it necessary to do a BURT restore? I tried doing one for c1susaux today and there were some errors for ASC channels, but the alignment sliders went right back to the proper place after reboot yesterday.

Quote:

I believe that the UR and LR magnets are stuck. There was no earth quake at 16:18 yesterday. Something had to kick it into this position. See 4days plot

Please advise  freeing details

 

  12493   Wed Sep 14 19:41:23 2016 JohannesUpdateGeneralPSL back to high power

Today's summary:

  • Replaced mirror in MC REFL path with R=10% BS and aligned beam on PD while still at low power
  • Replaced HR mirror in Transmon path at EY table with 50/50 BS. Alignment onto QPD not yet confirmed because we need IR from the YARM for it.
  • Put ND filters back on Transmon QPDs at both X and Y ends. For now I put all the filters on, for a combined OD of 1.6 at both ends (1.0 + 0.6 at YEND and 1.0 + 0.4 + 0.2 at XEND).
  • Put ND filter back on Transmon CCD on EY table.
  • Reverted MC autolocker to nominal, high power version.
  • Raised PSL output power back to nominal level by turning the waveplate. At the PSL shutter I measured a power of 1.03W. It occured to me too late that I realigned the PMC only afterwards and increased its transmission by a few percent, so I'll have to re-measure the actual PSL power.
  • MC is locked with its transmission back up to ~15,400 counts. The autolocker is not very good at obtaining the lock, as it seems to try to turn the VCO gain up too far and loses lock. The script probably needs a revision.
  • The YARM was pretty badly aligned. We used the green light to visually center the beam on the test masses AND had to go exploring with the TTs to see IR flashes in the first place. We got the YARM to lock to IR and were able to run the dither alignment. The maximum transmission we saw was on the order of 0.85. However, something strange is happening with the LSC control of the armlength. When the lock is engaged it drives PIT and YAW, which manifests itself in the OpLev signal and variable transmitted power on the TRY PDs. Osamu helped us diagnose this and was able to reduce the effect by tuning the POS gains to the individual ETMY OSEMS. The problem persisted even after using the new matrix diagonalization coefficients, we'll have to investigate this further and also take a look at the filters in the feedback path.
  • ITMX is still stuck and way out of alignment, so we couldn't even start with the green light in the XARM.
  12492   Wed Sep 14 09:44:09 2016 SteveUpdateVACpumpdown 79 completed

The pumpdown started at 9-12-2016

The IFO is at 5e-5 Torr vacuum normal after 73 day at atm.

ITMX needed to be freed and ETMY-UL is still misbehaving occasionally.

New pirani   and cold cathode gauges added at this vent. They were baked at 100 C for 6 hrs under vacuum.

Go to the Vac Rack to read IFO pressure from the gauge itself when Vac. Monitor is blank as it is now !

 

Quote:

P1 IFO pressure is 1 mTorr, valve configuration: vacuum normal, annulosses are pumped, RGA is not pumped.

THANKS to Chris !

The shutters can be opened with high power.

No communication error message still exist.

I will reboot c1vac1 and c1vac2 to get gauge communication with medm tomorrow.

 

 

 

Attachment 1: pd79.png
pd79.png
Attachment 2: pd79d2.jpg
pd79d2.jpg
Attachment 3: CCcg.jpg
CCcg.jpg
  12491   Wed Sep 14 09:33:54 2016 SteveUpdateGeneralITMX magnets are stucked

I believe that the UR and LR magnets are stuck. There was no earth quake at 16:18 yesterday. Something had to kick it into this position. See 4days plot

Please advise  freeing details.

Quote:

[Lydia,Teng]

Something strange happened to the ITMX osem reading around 4.pm. PDT as shown below.

Also the there was no response of the reading as we adjusted the PITCH and YAW. :(

Note that we restarted the slow machine: c1susaux,c1ausex this afternoon because of the unresponced interface.

 

 

 

Attachment 1: ITMXmagStuck.png
ITMXmagStuck.png
  12490   Tue Sep 13 19:18:43 2016 LydiaUpdateSUSDiagonalization in air

[Lydia, Teng]

We continued to work on the diagonalization scripts today and devised a way of choosing starting parameters that seems to work much better, and is easier to use, than tuning up to 15 parameters by hand per optic.

  • As before, the spectrum for each dof is estimated by using the "ideal" input matrix.
  • The starting guess for the peak frequency for each dof is the bin which achieves the maximum value of the spectrum between 0.4 and 1.5 Hz.
  • If another dof has a higher value at that frequency, the next highest peak is used. (Sometimes, for example, the peak in PIT at the POS frequency is stronger than the real POS peak!)
  • The peak height is initially guessed to be the spectrum value at the initial frequncy guess.
  • The width paramter Q can still be read from a file, but for all the times we tried, the peaks were found successfully if Q was initially guessed to be 300, so there might be no need to do this.
  • Spectra should still be examined to make sure the results make sense, and once we look at free swinging data in vacuum, we should compare the frequency results to the wiki values.
  • Reasonably good matrix values are saved to peakFit/inMats/1157630417. We got good diagonalization results for all but ITMY (see below). The values used for damping have not been overwritten.

We still noticed phase problems with ITMY, which appear to be preventing good diagonalization (See Attachment 1). Almost every degree of freedom has a significant imaginary part in the sensing matrix. We looked at the phases of the cross spectra in DDT and saw that indeed, the OSEM signals do not have the appropriate relative phases at the peak frequencies, especially in PIT and YAW (see Attachment 2: the phase at the peak is about 30 degrees when it should be 180). These phases are different for data takes ~24 hours apart, but are still wrong. We also looked at this information for ETMY and saw the correct behavior. We temporarily moved the pitch and yaw sliders for ITMY and looked at the OSEM response on a striptool, and the signals moved in the expected way. Can anyone suggest a reason why this would be happening? Is there another stretch of data (besides this past weekend) which would be good to compare to?

 

Attachment 1: ITMY_diag.jpg
ITMY_diag.jpg
Attachment 2: 38.png
38.png
  12489   Tue Sep 13 19:02:56 2016 TengUpdateGeneralITMX sensor

[Lydia,Teng]

Something strange happened to the ITMX osem reading around 4.pm. PDT as shown below.

Also the there was no response of the reading as we adjusted the PITCH and YAW. :(

Note that we restarted the slow machine: c1susaux,c1ausex this afternoon because of the unresponced interface.

 

 

Attachment 1: 47.png
47.png
Attachment 2: 34.png
34.png
  12488   Tue Sep 13 17:47:04 2016 SteveUpdateVACpumpdown 79 completed

P1 IFO pressure is 1 mTorr, valve configuration: vacuum normal, annulosses are pumped, RGA is off, not pumped.

THANKS to Chris !

The shutters can be opened with high power.

No communication error message still exist.

I will reboot c1vac1 and c1vac2 to get gauge communication with medm tomorrow.

 

Quote:

(Steve, Chris)

The pumpdown had stalled because of some ancient vacuum interlock code that prevented opening the valve V1 between the turbo pump and the main volume.

This interlock [0] compares the channels C1:Vac-P1_pressure and C1:Vac-PTP1_pressure, neither of which is functioning at the moment. The P1 channel apparently stopped reading sometime during the vent, and contained a value of ~700 torr, while the PTP1 channel contained 0. So the interlock code saw this huge apparent pressure difference and refused to move the valve.

To bypass this check, we used caput to enter a pressure of 0 for P1.

[0] /cvs/cds/caltech/state/from_luna/VacInterlock.st

 

ELOG V3.1.3-